OpenStudy (anonymous):

Express as a single natural log: ln(x-4)+ln(x+4)

OpenStudy (anonymous):

addition is multiplication and subtraction is division...

OpenStudy (anonymous):

\[ln(a)+ln(b)=ln(ab)\] \[ln(a)-ln(b)=\frac{ln(a)}{ln(b)}\]

OpenStudy (anonymous):

that's a lot of work, outkast lol... I prefer to have them look back at their notes, lectures, and textbook... reading is essential even in math

OpenStudy (anonymous):

thanks you guys. i got it now.

OpenStudy (anonymous):

takes no time at all if you type it in rather than using the editor